2x+3y=100----1
4x+5y=50----2

Get X & Y...


Answer by Alan3354(69443)

2x+3y=100 --> 4x+6y=200
4x+5y=50
4x+6y=200 Subtract
---------
0x- y=-150
y = 150
-------------
2x+3*150 = 100
2x = -350
x = -175